• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z. Chai

    Log entry time 00:00:38 on June29,2000

    Entry number 45126

    keyword=Swing shift summary


    16:00 swing shift starts
    16:18 Beam tripped due to IOC problem
    16:25 Beam back to Hall A(67uA)
    17:40 DAQ is having some problem. See halog 45102.
    Replayed the data, everything looks fine except that there is a noisy area
    in the FPP rx1 plane.
    22:35 End run 2613. Start run 2614. Prestart failed. Roc14 probelm. Kill coda,
    restart runcontrol doesn't work. Reset roc14, restart runcontrol.
    22:55 DAQ is working now.
    Start run 2617, dhist for e-arm is not working. Try to restart it, it reports
    "the et server is down". End run 2617, kill coda, start et sever and runcontrol.
    23:15 Start run 2618.

    Time Run Events Charge Comment
    16:00 2606 0.16M 23mC Beam trip because of IOC problem
    16:28 2607 1.3M 186mC
    17:17 2608 1.27M 183mC
    18:09 2609 1.3M 187mC
    18:58 2610 1.3M 187mC
    19:54 2611 1.3M 187mC
    20:44 2612 1.3 M 188mC
    21:34 2613 1.3M 187mC
    22:35 2614-2616 DAQ problem, Junk.
    22:58 2617 0.25M 36mC ET server was down.
    23:15 2618